OpenStudy (anonymous):

Which equation is in slope-intercept form and represents a line with slope 3 through the point (9, –4)?

OpenStudy (campbell_st):

well the equation of the line is y = 3x + b to find b, subsitute x = 9 and y = -4 and solve for b

OpenStudy (anonymous):

b=31?

OpenStudy (jdoe0001):

use the point-slope intercept form \(\large \begin{array}{llll} &x_1&y_1& \qquad slope = m = 3\\ (&9, &-4)\\\quad \\\quad \\ &&&\Large y-y_1=m(x-x_1) \end{array}\)

OpenStudy (campbell_st):

nearly, just check the sign of b... the value is correct

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."